Tim has 40 dimes and quarters with a total value of $7.75. How many dimes and how many quarters does he have?
Found 2 solutions by ikleyn, mananth:
Answer by ikleyn(52776)

.
Read the lesson
    - Three methods for solving standard (typical) coin word problems
in this site and learn from there how to solve such problems once for all.


Answer by mananth(16946)


x dimes'
y quarters

Equation for total coins

x + y = 40 .............1

Equation for amount

10 x + 25 y = 775.00 .............2
Eliminate y
multiply (1)by -25.00
Multiply (2) by 1
-25.00 x -25.00 y = -1000.00
10.00 25.00 y = 775.00
Add the two equations
-15.00 x = -225.00
/ -15.00
x = 15.00
plug value of x in (1)
1.00 x + 1.00 y = 40
15.00 + 1.00 y = 40
1.00 y = 25
y= 25
15 dimes' 25 quarters